Two equal positive charges q1=q2=2.0C are located at x=0, y=0.30m and x=0, y= - 0.30m, respectively. What are the magnitude and direction of the total electric force that q1 and q2 exert on the third charge Q=4.0C at x=0.40m, y=0?
1
Expert's answer
2019-11-18T10:38:01-0500



Define r1

"r_1=\sqrt{0.3^2+0.4^2}=\sqrt{0.09+0.16}=\sqrt{0.25}=0.5(m)"

Define r2

"r_2=\sqrt{0.3^2+0.4^2}=\sqrt{0.09+0.16}=\sqrt{0.25}=0.5(m)"

Define the angle "\alpha"

"\alpha=\arccos({0.4/0.5})=36,86^0"

Define Strength "F_1"

"F_1=\frac{1}{4\pi\epsilon_0}\cdot\frac{g_1Q}{r_1^2}=\frac{1}{4\pi\cdot8,85\cdot 10^{-12}(F/m)}\cdot\frac{2(C)\cdot 4(C)}{0.5^2(m^2)}=2,877\cdot10^{11}(N)"

from tog "q_1=q_2" and "r_1=r_2" follows that "F_2=F_1"

then

"F=2\cdot F_1 \cdot \cos(\alpha)=2\cdot 2,877 \cdot 10^{11}\cdot \cos(36,86^0)=4,603 \cdot 10^{11}(N)"